The units digit of a two-digit number is 2 more than the tens digit. If the number is divided by the sum of it's digits, the partial quotient is 4 and the remainder is 3. Find the number. I have figured out the number is 35, but again I am lost in finding the algebraic way to write it out.
Answer by AnlytcPhil(1810)

t = the tens digit
u = the units digit
10t+u = the number
t+u = sum of digits

The units digit of a two-digit number is 2 more than the tens digit.
So   u = t+2

If the number is divided by the sum of its digits, the partial quotient is 4 and the remainder is 3.
We divide as below.  Multiply the partial quotient 4
by the divisor, t+u, getting 4t+4u.  Then subtract that
from the dividend, 10t+u, and get remainder 6t-3u:

            4
   t+u)10t+ u
        4t+4u
        6t-3u = Remainder

And we are given that the remainder is 3. So

6t-3u = 3

So we have to solve this system of equations:

    u = t+2
6t-3u = 3

Solve that by substitution and get t = 3 and u = 5.

So the number is 35.

Edwin